rxcklesskaisher:

A company makes triangular plates for individual slices of pizza. For each plate, the base is 5 inches and the height is 16 inches. The area of the top of the plate is inches squared.

rxcklesskaisher:

20?

QuestionCoveBot:

Why would you think 20?

rxcklesskaisher:

idk i am getting no hrelp

jhonyy9:

on your past posted problem we said formula for area of a triangle is base time height divide 2 A = b×h/2 A = 5 ×16/2 = ?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rxcklesskaisher 20? \(\color{#0cbb34}{\text{End of Quote}}\) how you get this 20 ?